### v2

@@ Abstract
Context:Submillimeter/millimeter-wave single-dish telescopes offer two key advantages compared to interferometers: they can efficiently map larger portions of the sky, and they can recover larger spatial scales. Nonetheless, fluctuations in the atmosphere, the dominant noise source in ground-based observations, limit the accurate retrieval of signals from astronomical sources.
Aims:We introduce a versatile, user-friendly simulator to optimize scanning strategies and instrument designs to efficiently reduce atmospheric noise and filtering effects. We further use this tool to produce synthetic time streams and maps from hydrodynamical simulations, enabling a fair comparison between theory and reality.
Methods:To generate synthetic time-ordered data, we developed a multi-purpose telescope simulator called maria, which implements a suite of telescope and instrument designs intended to mimic current and future facilities. Each mock observatory scans through the atmosphere in a configurable pattern over the celestial object. We generate evolving and location-and-time-specific weather for each of the fiducial sites using a combination of satellite and ground-based measurements. While maria is a generic virtual telescope, this study specifically focuses on mimicking broadband bolometers observing at GHz.
Results: To validate our virtual telescope, we compare the mock time streams with real MUSTANG-2 observations and find that they are quantitatively similar by conducting a -sample Anderson-Darling test resulting in a value of . Subsequently, we image the time-ordered data to create noise maps and realistic mock observations of clusters of galaxies for both MUSTANG-2 and an instrument concept for the 50 m Atacama Large Aperture Submillimeter Telescope (AtLAST). Furthermore, using maria, we find that a 50 m dish provides the highest levels of correlation of atmospheric signals across adjacent detectors compared to smaller apertures (e.g., 42-cm and 6-m survey experiments), facilitating removal of atmospheric signal on large scales.
Conclusions:We demonstrated the usage of maria for facilitating detailed forecasts for frontier instruments, maximizing their scientific output.
maria: A novel simulator for forecasting (sub-)mm observations
Key Words.:
@@
Astronomical instrumentation, methods and techniques – Atmospheric effects – Techniques: image processing – Submillimeter: general – Galaxies: clusters: intracluster medium1 Introduction
@@
The most sensitive (sub-)millimeter [hereafter, (sub-)mm] user facility yet built is the Atacama Large Millimeter/submillimeter Array (ALMA), which outperforms current single-dish telescopes and other interferometers in terms of point source sensitivity due to its large collecting area and excellent site. However, interferometers like ALMA are optimized to achieve sub-arcsecond resolution in the (sub-)mm, but this comes at the expense of being unable to constrain large spatial scales, which are probed by short antenna spacings. This in turn is simply due to there being a minimum distance between two antennas that is needed to avoid collisions while tracking a source. This limit, known as the shadowing limit, prevents zero-spacing information and results in a biased recovery of the total integrated flux. Previous works have shown that even at scales the size of the synthesized beam (i.e., the interferometric resolution element), only a fraction of the true flux is recovered (see, e.g., Plunkett et al., 2023). Simulations indicate that to address the loss of information at larger angular scales and to provide good overlap in the Fourier domain, one should rely on single-dish telescopes with a dish size of at least three times the size of the interferometric array element. This is necessary to provide the complete Fourier sampling required for high spatial dynamic range and unbiased image reconstruction (Frayer, 2017; Plunkett et al., 2023).
@@
On the other hand, large single-dish telescopes have their own observational limitations. For example, due to the faintness of the signal compared to the Earth’s atmospheric emission, drifts in detector gain and response, and the fact most bolometer-based instruments are not set up to perform an absolute temperature calibration, bolometric arrays require differential measurements (e.g., going on and off source) to obtain an unbiased flux estimate. Single-dish user-facility telescopes like the Green Bank Telescope (GBT) therefore adopt a Lissajous daisy scanning pattern (e.g. Romero et al., 2020). These scanning trajectories impact the observed sky and introduce filtering effects, such as a higher sensitivity in the center compared to the outskirts.
@@
The current implementation of maria is flexible and can be customized to various types of instruments, such as heterodyne and integrated field units, to generate mock datacubes with spectral information. However, in this study specifically, we limit maria to generate synthetic broadband bolometric observations. In particular, we restrict maria three-fold: (i) We mimic only two facilities, of which one already exists (the GBT with the MUSTANG-2 instrument) which allows testing the maria output with real observations, and a future facility, namely the 50 meter Atacama Large Aperture Submillimeter Telescope (AtLAST; Ramasawmy et al. 2022; Mroczkowski et al. 2023) with a possible bolometer instrument concept.222We note that AtLAST will host a multitude of different instruments, also including spectroscopic receivers. The tool itself includes configuration files to easily simulate more facilities, such as the Large Millimeter Telescope (Kaercher & Baars, 2000, LMT;), as well as current CMB survey experiments such as the South Pole Telescope (Carlstrom et al., 2011; Bleem et al., 2015, 2020, SPT;) and the Atacama Cosmology Telescope (ACT; Swetz et al., 2011; Thornton et al., 2016; Henderson et al., 2016), imminent ones like Simons Observatory (SO, Ade et al., 2019) and CCAT-prime (Fred Young Submm Telescope or FYST; CCAT-Prime Collaboration 2023), and future ones like CMB-S4 (Abazajian et al., 2016). (ii) We focus on the GHz regime, which is the high-frequency regime for the GBT, and the low-frequency regime for a telescope such as AtLAST. We note that maria can also simulate the atmosphere at higher frequencies up to where the atmosphere is still transmissive well into the submm. (iii) Finally, even though maria is generic and can be employed for a wide variety of science cases, we focus here on a specific one: the intracluster medium seen via the thermal Sunyaev-Zeldovich (SZ; Sunyaev & Zeldovich 1970, 1972) effect. Briefly, the thermal SZ effect is a spectral distortion to the CMB, presents as a faint, extended decrement at frequencies GHz, and is a probe of gas thermal energy (see, e.g., Mroczkowski et al. 2019 for a review, as well as Di Mascolo et al. in prep for a more comprehensive overview of key AtLAST SZ science cases).
@@ 2 A generic virtual single dish telescope
The simulator we built is a continuation of the atmospheric modeling work done by Morris et al. (2022, 2024 in prep). Therefore, we refer to those works and references therein for an in-depth discussion of atmospheric modeling. Here, we briefly describe how the maria simulation works. The simulator consists of five main components: 1) a telescope and instrument design, 2) a scanning strategy, 3) an input astronomical map, 4) a site and time location to generate atmospheric weather conditions, and 5) additional additive noise and detector inefficiencies. These components are visualized in Figure 1 and will be discussed separately in the following five subsections. For the full overview of all the parameters and corresponding tutorials, we refer to the maria website.1
With this setup, we can directly link fundamental design choices, such as primary size and the number of detectors, to a scientific output. Thus, we established a tool for communication between instrument scientists and observers, which is crucial for the development of future facilities.
@@ 2.1 Instrument configurations & Telescope designs
In maria, each simulation is characterized by a set of fundamental parameters regarding both the telescope and instrument design. Specifically, we use the primary size, the field of view (FoV), and the azimuthal and elevation bounds of the telescope for modeling the simulations. Further, maria positions detectors in the array evenly to fill the FoV in a hexagonally packed configuration. In the current implementation, each detector is specified with a default frequency band that has a flat response function (e.g., a flat passband from 140-160 GHz, centered at 150 GHz). maria also hosts the ability to import passbands.
@@ 2.1.2 AtLAST
AtLAST (Klaassen et al., 2020; Ramasawmy et al., 2022) is a concept for a next-generation high-throughput submm telescope to be sited high in the Atacama Desert of Chile, on Llano de Chajnantor (i.e., the same plateau as that on which ALMA resides) at an elevation of approximately 5100 m above sea level. AtLAST is currently undergoing a design study funded by the European Commission.333See https://cordis.europa.eu/project/id/951815. Here, we define throughput, often referred to as étendue, as the collecting area of the primary mirror times the telescope FoV. With a surface accuracy of 20 m, AtLAST will be capable of observations up to 950 GHz (i.e., in the 350 m atmospheric window accessible in top octile conditions there). AtLAST will include dedicated instrument spaces for six extremely large receivers, two of which can completely access the full 4.6 meter focal plane corresponding to the 2 FoV (see Mroczkowski et al., 2023).
@@ 2.3 Input astronomical map
Often in the millimeter-wave regime, astronomical signals are contaminated by fore- and background sources, like the CMB, the Galaxy, and the Cosmic Infrared Background (CIB), where the latter is generally the unresolved confusion-limited submm/FIR emission dominated by dusty galaxies from different epochs, peaking between (see, i.e., Devlin et al., 2009; MacCrann et al., 2023; Madhavacheril et al., 2023). In future implementations of maria, there will be an option to include these components through a power spectra modelization, which will be jointly mapped with the atmosphere. However, for the current version, we recommend manually adding these contamination sources to the input file when making forecasts for survey experiments, such as detecting the primary CMB or when performing line intensity mapping experiments.
@@ 2.4 Atmospheric modeling
The atmospheric modeling in maria is the backbone of the corruption model of the simulation. Therefore, this subsection details how the atmospheric modeling is set up and implemented. In particular, we discuss which weather data is used to generate local-and-time specific initial conditions (Section 2.4.1), how we generate atmospheric emission (Section 2.4.2) and corresponding turbulence from the initial conditions (Section 2.4.3), how we let the atmosphere evolve with time while scanning over it (Section 2.4.4), and how all this is implemented in maria (Section 2.4.5). For a more comprehensive overview of how the atmospheric model is built and how it compares with ACT observations, we refer to works of Morris et al. (2022, 2024 in prep).
@@ 2.4.1 Weather data
To simulate weather conditions at astronomical sites, maria uses reanalysis data from the European Centre for Medium-Range Weather Forecasts (ECMWF, Hersbach et al., 2020, where the fifth-generation atmospheric reanalysis is known as ERA5).444See https://thomaswmorris.com/maria/sites for a list of supported sites. Reanalysis data is a mixture of weather observations with past short-range forecasts rerun with modern weather forecasting models from Hersbach et al. (2020). It further has the advantage of providing a characterization of the entire atmospheric column, which can often be quite different from ground-based weather measurements.
@@ 2.4.2 Atmospheric emission and opacity
As the dominant source of fluctuation in atmospheric emission and opacity for submillimeter telescopes is precipitable water vapor (PWV), we simulate atmospheric fluctuations by perturbing the line-of-sight water vapor of the receiver. We do this by generating a set of layers of turbulent fluctuations and scaling them appropriately according to site-specific weather data derived from ERA5. Typical minute-to-minute variations in PWV are on the order of a few percent of the total (Morris et al., 2024 in prep), though the relative fluctuation is a variable parameter in maria. We compute the atmospheric emission and opacity spectra using the vertical profiles of temperature, humidity, ozone, and wind velocity using the am atmospheric model (Paine, 2018) for a given site, altitude, and observing elevation. Samples of these spectra for different levels of PWV at the Green Bank site are shown in Figure 3.
@@ 2.4.3 Modeling turbulence
Atmospheric emission comprises a large, constant brightness from the atmosphere (which is almost always subtracted out of the timestream) but also a smaller time-dependent brightness, due mostly to fluctuations atmosphere temperature and water vapor density. Due to turbulent mixing between atmospheric layers of different temperature and humidity, local turbulent variations introduce spatial and temporal fluctuations in the atmosphere; see Tatarski 1961 for a statistical description of this phenomenon. For small angular separations, Morris et al. (2022) showed that the angular covariance of turbulent fluctuations in the line-of-sight atmospheric emission can be modeled as:
where is the angular separation on the sky, is the distance from the observer, is the modified Bessel function of the second kind of order , and is the scaling of atmospheric brightness fluctuations as a function of distance from the observer, and is the turbulent outer scale.555An analogous relation holds for opacity fluctuations. This equation can be interpreted as a series of “integrated” layers of turbulence with a modified spectrum. For small angular separations, we recover a power law in the structure function which arises from:
where is the turbulent covariance function in Eq. (2). This power law has been observed in, among others, Consortini & Ronchi (1972), Conan et al. (2000), Tokovinin (2002), and Morris et al. (2022). We use the value of meters for the turbulent outer scale , which comes from measurements done by Errard et al. (2015) and Morris et al. (2024 in prep). We further assume a “frozen-in” model for the turbulence (Lay, 1997), meaning that fluctuations are only induced by the motion along the line-of-sight with respect to the atmosphere as translated by the wind. For each detector , we can compute its atmosphere-relative coordinates as
@@ 2.4.4 Modeling the evolving atmosphere
The maria code simulates turbulence using an auto-regressive version of the covariance matrix method (see e.g., Assémat et al. 2006). This method is preferred to Fourier-based methods (see e.g., Jia et al. 2015) for the simulations in this work, as they employ scans over a small area on the sky with long integration times. This means that the set of angular positions as described by Eq. 4 will comprise a long, thin strip of atmosphere which is much more efficiently generated with the former method. The covariance matrix method assumes that perturbations in atmosphere temperature and PWV due to turbulent mixing are described by a stationary Gaussian process with some covariance matrix. Generating a realization of such a process requires inverting the covariance matrix, which scales with cubic complexity in the number of points and rapidly becomes computationally infeasible for . We can instead generate a realization iteratively by conditioning a posterior on a sparse sampling of already-realized points and sampling from it. Considerations of sampling methods for balancing performance and realism are addressed in Morris et al. in prep., along with the trade-offs of different methods of turbulent generation. We compute the angular covariance as described by Eq. 2, and adjust the generated turbulent distribution to recover the mean and variance matching the simulation parameters.

### v3

@@ Abstract
Submillimeter/millimeter-wave single-dish telescopes offer two key advantages compared to interferometers: they can efficiently map larger portions of the sky, and they can recover larger spatial scales. Nonetheless, fluctuations in the atmosphere, the dominant noise source in ground-based observations, limit the accurate retrieval of signals from astronomical sources. We introduce a versatile, user-friendly simulator to optimize scanning strategies and instrument designs to efficiently reduce atmospheric noise and filtering effects. We further use this tool to produce synthetic time streams and maps from hydrodynamical simulations, enabling a fair comparison between theory and reality. To generate synthetic time-ordered data, we developed a multi-purpose telescope simulator called maria, which implements a suite of telescope and instrument designs intended to mimic current and future facilities. Each mock observatory scans through the atmosphere in a configurable pattern over the celestial object. We generate evolving and location-and-time-specific weather for each of the fiducial sites using a combination of satellite and ground-based measurements. While maria is a generic virtual telescope, this study specifically focuses on mimicking broadband bolometers observing at GHz. To validate our virtual telescope, we compare the mock time streams with real MUSTANG-2 observations and find that they are quantitatively similar by conducting a -sample Anderson-Darling test resulting in a value of . Subsequently, we image the time-ordered data to create noise maps and realistic mock observations of clusters of galaxies for both MUSTANG-2 and an instrument concept for the 50 m Atacama Large Aperture Submillimeter Telescope (AtLAST). Furthermore, using maria, we find that a 50 m dish provides the highest levels of correlation of atmospheric signals across adjacent detectors compared to smaller apertures (e.g., 42-cm and 6-m survey experiments), facilitating removal of atmospheric signal on large scales. We demonstrated the usage of maria for facilitating detailed forecasts for frontier instruments, maximizing their scientific output.
keywords:
@@
Astronomical instrumentation, methods and techniques – Atmospheric effects – Galaxy Clusters1 Introduction
@@
The most sensitive (sub-)millimeter [hereafter, (sub-)mm] user facility yet built is the Atacama Large Millimeter/submillimeter Array (ALMA), which outperforms current single-dish telescopes and other interferometers in terms of point source sensitivity due to its large collecting area and excellent site. However, interferometers like ALMA are optimized to achieve sub-arcsecond resolution in the (sub-)mm, but this comes at the expense of being unable to constrain large spatial scales, which are probed by short antenna spacings. This is simply due to the minimum distance between two antennas needed to avoid collisions while tracking a source. This limit, known as the shadowing limit, prevents zero-spacing information and results in a biased recovery of the total integrated flux. Previous works have shown that even at scales the size of the synthesized beam (i.e., the interferometric resolution element), only a fraction of the true flux is recovered (Plunkett et al., 2023; Bonanomi et al., 2024). Simulations indicate that to address the loss of information at larger angular scales and to provide good overlap in the Fourier domain, one should rely on single-dish telescopes with a dish size of at least three times the size of the interferometric array element. This is necessary to provide the complete Fourier sampling required for high spatial dynamic range and unbiased image reconstruction (Frayer, 2017; Plunkett et al., 2023).
@@
On the other hand, large single-dish telescopes have their own observational limitations. For example, due to the faintness of the signal compared to the Earth’s atmospheric emission, drifts in detector gain and response, and the fact that most bolometer-based instruments are not set up to perform an absolute temperature calibration, bolometric arrays require differential measurements (e.g., going on and off source) to obtain an unbiased flux estimate. Single-dish user-facility telescopes like the Green Bank Telescope (GBT) therefore adopt a Lissajous daisy scanning pattern (e.g. Romero et al., 2020). These scanning trajectories impact the observed sky and introduce filtering effects, such as a higher sensitivity in the center compared to the outskirts of the observation.
@@
The current implementation of maria is flexible and can be customized to various types of instruments, such as heterodyne and integrated field units, to generate mock datacubes with spectral information. However, in this study specifically, we limit maria to generate synthetic broadband bolometric observations. In particular, we restrict maria three-fold: (i) We mimic only two facilities, of which one already exists (the GBT with the MUSTANG-2 instrument) which allows testing the maria output with real observations, and a future facility, namely the 50 meter Atacama Large Aperture Submillimeter Telescope (AtLAST; Ramasawmy et al. 2022; Mroczkowski et al. 2023; Booth et al. 2024; Mroczkowski et al. 2024; Reichert et al. 2024)222See also: https://atlast-telescope.org with a possible bolometer instrument concept.333We note that AtLAST will host a multitude of different instruments, also including spectroscopic receivers. The tool itself includes configuration files to easily simulate more facilities, such as the Large Millimeter Telescope (Kaercher & Baars, 2000, LMT;), as well as current CMB survey experiments such as the South Pole Telescope (Carlstrom et al., 2011; Bleem et al., 2015, 2020, SPT;) and the Atacama Cosmology Telescope (ACT; Swetz et al., 2011; Thornton et al., 2016; Henderson et al., 2016), imminent ones like Simons Observatory (SO, Ade et al., 2019) and CCAT-prime (Fred Young Sub-mm Telescope or FYST; CCAT-Prime Collaboration 2023), and future ones like CMB-S4 (Abazajian et al., 2016). (ii) We focus on the GHz regime, which is the high-frequency regime for the GBT, and the low-frequency regime for a telescope such as AtLAST. We note that maria can also simulate the atmosphere at higher frequencies up to where the atmosphere is still transmissive well into the sub-mm. (iii) Finally, even though maria is generic and can be employed for a wide variety of science cases, we focus here on a specific one: the intracluster medium seen via the thermal Sunyaev-Zeldovich (SZ; Sunyaev & Zeldovich 1970, 1972) effect. Briefly, the thermal SZ effect is a spectral distortion to the CMB, presents as a faint, extended decrement at frequencies GHz, and is a probe of gas thermal energy (see, e.g., Mroczkowski et al. 2019 for a review, as well as Di Mascolo et al. 2024 for a more comprehensive overview of key AtLAST SZ science cases).
@@ 2 A generic virtual single dish telescope
The simulator we built is a continuation of the atmospheric modeling work done by Morris et al. (2022, 2024). Therefore, we refer to those works and references therein for an in-depth discussion of atmospheric modeling. Here, we briefly describe how the maria simulation works. The simulator consists of five main components: 1) a telescope and instrument design, 2) a scanning strategy, 3) an input astronomical map, 4) a site and time location to generate atmospheric weather conditions, and 5) additional additive noise and detector inefficiencies. These components are visualized in Figure 1 and will be discussed separately in the following five subsections. For the full overview of all the parameters and corresponding tutorials, we refer to the maria documentation.1 With this setup, we can directly link fundamental design choices, such as primary size and the number of detectors, to a scientific output. Thus, we established a tool for communication between instrument scientists and observers, which is crucial for the development of future facilities.
@@ 2.1 Instrument configurations & Telescope designs
In maria, each simulation is characterized by a set of fundamental parameters regarding both the telescope and instrument design. Specifically, we use the primary size, the field of view (FoV), the number of detectors, and the azimuthal and elevation bounds of the telescope for modeling the simulations. Further, maria positions detectors in the array evenly to fill the FoV in a hexagonally packed configuration. In the current implementation, each detector is specified with a default frequency band with a flat response function (e.g., a flat passband from 140-160 GHz, centered at 150 GHz). maria also hosts the ability to import passbands.
@@ 2.1.2 AtLAST
AtLAST (Klaassen et al., 2020; Ramasawmy et al., 2022) is a concept for a next-generation high-throughput sub-mm telescope to be sited high in the Atacama Desert of Chile, on Llano de Chajnantor at an elevation of approximately 5100 m above sea level (i.e., the same plateau as that on which ALMA resides). AtLAST underwent a design study, funded by the European Commission,444See https://cordis.europa.eu/project/id/951815. and has been selected for a new infrastructure development grant, also funded by the European Commission, starting in January 2025. Among other activities, AtLAST will continue to improve its technology readiness levels and perform scientific forecasts using, for example, the maria code presented here. With a surface accuracy of 20 m, AtLAST will be capable of observations up to 950 GHz (i.e., in the 350 m atmospheric window accessible in top octile conditions there). AtLAST will include dedicated instrument spaces for six extremely large receivers, two of which can completely access the full 4.7-meter focal plane corresponding to the 2∘ FoV (see Gallardo et al., 2024; Kiselev et al., 2024; Puddu et al., 2024; Reichert et al., 2024; Mroczkowski et al., 2024).
@@ 2.3 Input astronomical map
Often in the millimeter-wave regime, astronomical signals are contaminated by fore- and background sources, like the CMB, the Galaxy, and the Cosmic Infrared Background (CIB), where the latter is generally the unresolved confusion-limited sub-mm/FIR emission dominated by dusty galaxies from different epochs, peaking between (see, i.e., Devlin et al., 2009; MacCrann et al., 2024; Madhavacheril et al., 2024). In future implementations of maria, there will be an option to include these components through a power spectra modelization, which will be jointly mapped with the atmosphere. However, for the current version, we recommend manually adding these contamination sources to the input file when making forecasts for survey experiments, such as detecting the primary CMB or when performing line intensity mapping experiments.
@@ 2.4 Atmospheric modeling
The atmospheric modeling in maria is the backbone of the corruption model of the simulation. Therefore, this subsection details how the atmospheric modeling is set up and implemented. In particular, we discuss which weather data is used to generate local-and-time specific initial conditions (Section 2.4.1), how we generate atmospheric emission (Section 2.4.2) and corresponding turbulence from the initial conditions (Section 2.4.3), how we let the atmosphere evolve with time while scanning over it (Section 2.4.4), and how all this is implemented in maria (Section 2.4.5). For a more comprehensive overview of how the atmospheric model is built and how it compares with ACT observations, we refer to works of Morris et al. (2022, 2024).
@@ 2.4.1 Weather data
To simulate weather conditions at astronomical sites, maria uses reanalysis data from the European Centre for Medium-Range Weather Forecasts (ECMWF, Hersbach et al., 2020, where the fifth-generation atmospheric reanalysis is known as ERA5).555See https://thomaswmorris.com/maria/sites for a list of supported sites. Reanalysis data is a mixture of weather observations with past short-range forecasts rerun with modern weather forecasting models from Hersbach et al. (2020). It further has the advantage of providing a characterization of the entire atmospheric column, which can often be quite different from ground-based weather measurements.
@@ 2.4.2 Atmospheric emission and opacity
As the dominant source of fluctuation in atmospheric emission and opacity for submillimeter telescopes is precipitable water vapor (PWV), we simulate atmospheric fluctuations by perturbing the line-of-sight water vapor of the receiver. We do this by generating a set of layers of turbulent fluctuations and scaling them appropriately according to site-specific weather data derived from ERA5. Typical minute-to-minute variations in PWV are on the order of a few percent of the total (Morris et al., 2024), though the relative fluctuation is a variable parameter in maria. We compute the atmospheric emission and opacity spectra using the vertical profiles of temperature, humidity, ozone, and wind velocity using the am atmospheric model (Paine, 2018) for a given site, altitude, and observing elevation. Samples of these spectra for different levels of PWV at the Green Bank site are shown in Figure 3.
@@ 2.4.3 Modeling turbulence
Atmospheric emission comprises a large, constant brightness from the atmosphere (which is almost always subtracted out of the timestream) but also a smaller time-dependent brightness due mostly to fluctuations in atmosphere temperature and water vapor density. Due to turbulent mixing between atmospheric layers of different temperature and humidity, local turbulent variations introduce spatial and temporal fluctuations in the atmosphere; see Tatarski 1961 for a statistical description of this phenomenon. For small angular separations, Morris et al. (2022) showed that the angular covariance of turbulent fluctuations in the line-of-sight atmospheric emission can be modeled as:
where is the angular separation on the sky, is the distance from the observer, is the modified Bessel function of the second kind of order , and is the scaling of atmospheric brightness fluctuations as a function of distance from the observer, and is the turbulent outer scale.666An analogous relation holds for opacity fluctuations. This equation can be interpreted as a series of “integrated” layers of turbulence with a modified spectrum. For small angular separations, we recover a power law in the structure function which arises from:
where is the turbulent covariance function in Eq. (2). This power law has been observed in, among others, Consortini & Ronchi (1972), Conan et al. (2000), Tokovinin (2002), and Morris et al. (2022). We use the value of meters for the turbulent outer scale , which comes from measurements done by Errard et al. (2015) and Morris et al. (2024). We further assume a “frozen-in” model for the turbulence (Lay, 1997), meaning that fluctuations are only induced by the motion along the line-of-sight with respect to the atmosphere as translated by the wind. For each detector , we can compute its atmosphere-relative coordinates as
@@ 2.4.4 Modeling the evolving atmosphere
The maria code simulates turbulence using an auto-regressive version of the covariance matrix method (see e.g., Assémat et al. 2006). This method is preferred to Fourier-based methods (see e.g., Jia et al. 2015) for the simulations in this work, as they employ scans over a small area on the sky with long integration times. This means that the set of angular positions as described by Eq. 4 will comprise a long, thin strip of atmosphere which is much more efficiently generated with the former method. The covariance matrix method assumes that perturbations in atmosphere temperature and PWV due to turbulent mixing are described by a stationary Gaussian process with some covariance matrix. Generating a realization of such a process requires inverting the covariance matrix, which scales with cubic complexity in the number of points and rapidly becomes computationally infeasible for . We can instead generate a realization iteratively by conditioning a posterior on a sparse sampling of already-realized points and sampling from it. Considerations of sampling methods for balancing performance and realism are addressed in Morris et al. (2024), along with the trade-offs of different methods of turbulent generation. We compute the angular covariance as described by Eq. 2, and adjust the generated turbulent distribution to recover the mean and variance matching the simulation parameters.